The next four questions are based on the following information: Coco Chips Company, a retail store, prepares its master budge Just 10. Cost of Merchandise Purchase.
Answer #1

9) Account receivable at May 31 = 548000*48% = 263040

So answer is b) $263040

10) Cost of merchandise purchase = (578000*60%)+(461000*60%*20%)-(578000*60%*20%) = $332760

So answer is d) $332760